# Differential Encoding for Improved Representation Learning over Graphs

### v1

@@ III-A Preliminaries
$\mathbf{h}_{u}^{(k)}$ $(\displaystyle=\text{Update}^{(k)}(\text{Aggregate}^{(k)}($1)
@@ III-C Differential Encoding in Aggregation
[…]
@@ III-D Model Architecture
$\mathbf{h\_local}_{u}^{(k)}$ $(\displaystyle=\text{Update}^{(k)}(\text{Aggregate}^{(k)}($8)
@@
[…]
@@
$=\text{Concat}(\mathbf{H}_{1}^{(k)},...,\mathbf{H}_{N_{h}}^{(k)})\mathbf{W}_{MHA}^{(k)},$
@@
[…]
@@ IV Experiments
GCN [28] 90.705$\pm$0.218 55.710$\pm$0.381
MoNet [29] 90.805$\pm$0.032 54.655$\pm$0.518
GraphSAGE [17] 97.312$\pm$0.097 65.767$\pm$0.308
GIN [30] 96.485$\pm$0.252 55.255$\pm$1.527
GCNII [1] 90.667$\pm$0.143 56.081$\pm$0.198
PNA [31] 97.94$\pm$0.12 70.35$\pm$0.63
DGN [32] 72.838$\pm$0.417
Cy2C-GNNs [33] 97.772$\pm$0.001 64.285$\pm$0.005
ARGNP [34] 73.90$\pm$0.15
CRaWl [35] 97.944$\pm$0.050 69.013$\pm$0.259
GIN-AK+ [36] 72.19$\pm$0.13
3WLGNN [37] 95.075$\pm$0.961 59.175$\pm$1.593
EGT [38] 98.173$\pm$0.087 68.702$\pm$0.409
GatedGCN [18] 97.340$\pm$0.143 67.312$\pm$0.311
GatedGCN + SSFG [7] 97.985$\pm$0.032 71.938$\pm$0.190
EdgeGCN [39] 98.432$\pm$0.059 76.127$\pm$0.402
Exphormer [3] 98.550$\pm$0.039 74.754$\pm$0.194
TIGT [40] 98.230$\pm$0.133 73.955$\pm$0.360
GRIT [41] 98.108$\pm$0.111 76.468$\pm$0.881
GPS [11] (base model) 98.051$\pm$0.126 72.298$\pm$0.356
Ours 98.558$\pm$0.057 79.067$\pm$0.269
@@
GCN [28] 0.1268$\pm$0.0060 0.0841$\pm$0.0010
GINE [42] 0.1265$\pm$0.0076 0.1339$\pm$0.0044
GCNII [1] 0.1698$\pm$0.0080 0.1404$\pm$0.0011
GatedGCN [18] 0.2873$\pm$0.0219 0.2641$\pm$0.0045
GatedGCN + RWSE [11] 0.2860$\pm$0.0085 0.2574$\pm$0.0034
Transformer + LapPE [43] 0.2694$\pm$0.0098 0.2618$\pm$0.0031
SAN + LapPE [43] 0.3230$\pm$0.0039 0.2592$\pm$0.0158
SAN + RWSE [43] 0.3216$\pm$0.0027 0.2434$\pm$0.0156
Exphormer [3] 0.3975$\pm$0.0037 0.3455$\pm$0.0009
GPS [11] 0.3748$\pm$0.0109 0.3412$\pm$0.0044
Ours 0.4242$\pm$0.0011 0.3567$\pm$0.0026
@@
GCN 0.1321$\pm$0.0007 0.3791$\pm$0.0004 0.8256$\pm$0.0006 0.3234$\pm$0.0006
GCNII 0.1325$\pm$0.0009 0.3607$\pm$0.0003 0.8116$\pm$0.0009 0.3161$\pm$0.0004
GINE 0.1337$\pm$0.0013 0.3642$\pm$0.0043 0.8147$\pm$0.0062 0.3180$\pm$0.0027
GatedGCN 0.1279$\pm$0.0018 0.3783$\pm$0.0004 0.8433$\pm$0.0011 0.3218$\pm$0.0011
GatedGCN+RWSE 0.1288$\pm$0.0013 0.3808$\pm$0.0006 0.8517$\pm$0.0005 0.3242$\pm$0.0008
Transformer+LapPE 0.1221$\pm$0.0011 0.3679$\pm$0.0033 0.8517$\pm$0.0039 0.3174$\pm$0.0020
SAN+LapPE 0.1355$\pm$0.0017 0.4004$\pm$0.0021 0.8478$\pm$0.0044 0.3350$\pm$0.0003
SAN+RWSE 0.1312±0.0016 0.4030$\pm$0.0008 0.8550$\pm$0.0024 0.3341$\pm$0.0006
Graph Diffuser [44] 0.1369$\pm$0.0012 0.4053$\pm$0.0011 0.8592$\pm$0.0007 0.3388$\pm$0.0011
GPS [11] 0.1471$\pm$0.0008 0.3937$\pm$0.0019 0.8526$\pm$0.0014 0.3337$\pm$0.0006
Ours 0.1490$\pm$0.0004 0.4134$\pm$0.0010 0.8557$\pm$0.0003 0.3459$\pm$0.0005
@@
GCN [28] 0.6839$\pm$0.0084
GCN + virtual node 0.6857$\pm$0.0061
GIN [30] 0.6892$\pm$0.0100
GIN + virtual node 0.7037$\pm$0.0107
DeeperGCN 0.7712$\pm$0.0071
ExpC [45] 0.7976$\pm$0.0072
K-Subtree SAT 0.7522$\pm$0.0056
GPS [11] 0.8015$\pm$0.0033
Ours 0.8096$\pm$0.0029
@@
GCN 0.5930$\pm$0.0023
GINE 0.5498$\pm$0.0079
GCNII [1] 0.5543$\pm$0.0078
GatedGCN 0.5864$\pm$0.0077
Gated + RWSE 0.6069$\pm$0.0035
Transformer+LapPE 0.6326$\pm$0.0126
SAN+LapPE 0.6384$\pm$0.0121
SAN+RWSE 0.6439$\pm$0.0075
Exphormer [3] 0.6527$\pm$0.0043
GPS [11] 0.6535$\pm$0.0041
Ours 0.6608$\pm$0.0046
@@
MPNN only 97.989$\pm$0.118 76.823$\pm$0.219 0.3579$\pm$0.0047 0.2446$\pm$0.0027
MPNN + differential encoding 98.285$\pm$0.080 77.820$\pm$0.091 0.3666$\pm$0.0035 0.2500$\pm$0.0012
Global attention only 97.564$\pm$0.146 45.237$\pm$1.349 0.2748$\pm$0.0052 0.2535$\pm$0.0047
Global attention + differential encoding 97.912$\pm$0.062 64.468$\pm$0.537 0.2995$\pm$0.0013 0.2608$\pm$0.0013
Full Model 98.528$\pm$0.057 79.067$\pm$0.269 0.4242$\pm$0.0011 0.3567$\pm$0.0026
@@
MPNN only 0.3306$\pm$0.0004 0.6131$\pm$0.0043 0.7995$\pm$0.0042
MPNN + differential encoding 0.3406$\pm$0.0002 0.6201$\pm$0.0025 0.8059$\pm$0.0023
Global attention only 0.3276$\pm$0.0012 0.6207$\pm$0.0088 0.0948$\pm$0.0000
Global attention + differential encoding 0.3277$\pm$0.0018 0.6355$\pm$0.0067 0.0948$\pm$0.0000
Full Model 0.3459$\pm$0.0005 0.6608$\pm$0.0046 0.8096$\pm$0.0029
@@ IV-B Experimental Results
GCN 90.120$\pm$0.145 54.142$\pm$0.394
GCN + differential enc. 94.728$\pm$0.137 62.020$\pm$0.362
(4.608$\uparrow$) (7.878$\uparrow$)
GAT 95.535$\pm$0.205 64.223$\pm$0.455
GAT + differential enc. 97.163$\pm$0.154 67.373$\pm$0.415
(1.628$\uparrow$) (3.120$\uparrow$)

### v2

@@ III-A Preliminaries
$\mathbf{h}_{u}^{(k)}$ $=\text{Update}^{(k)}(\text{Aggregate}^{(k)}($ (1)
@@ III-C Differential Encoding in Aggregation
$\hskip 20.0pt\text{diff\_enc}(\sum_{v\in\mathcal{A}}\mathbf{m}^{(k)}_{v\to u}-\mathbf{m}^{(k)}_{u\to u})$
@@ III-D Model Architecture
$\mathbf{h\_local}_{u}^{(k)}$ $=\text{Update}^{(k)}(\text{Aggregate}^{(k)}($ (8)
@@
$\hskip 50.0pt\text{diff\_enc}_{l}^{(k)}(\sum_{v\in\mathcal{N}{(u)}}\mathbf{m}_{v\to u}^{(k)}-\mathbf{m}_{u\to u}^{(k)})).$
@@
$=\text{Concat}(\mathbf{H}_{1}^{(k)},.,\mathbf{H}_{N_{h}}^{(k)})\mathbf{W}_{MHA}^{(k)},$
@@
$\hskip 20.0pt\text{diff\_enc}_{g,i}^{(k)}(\text{Softmax}(\frac{\mathbf{Q}_{i}^{(k)}{\mathbf{K}_{i}^{(k)}}^{T}}{\sqrt{d_{q}^{(k)}}})\mathbf{V}_{i}^{(k)}$-
@@ IV Experiments
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV-B Experimental Results
[…]